The resistance of the primary and secondary coils
Answer #1

In transformer i P / i S = N s / Np

We know i P / i S = R s / R P

Since current is inversely proportional to resistsnce.

From above two equations , N s / Np =R s / R P

                                                    = 12 ohm / 71 ohm

                                                    = 12 /71